Multiple Product Break-Even and Net Income Planning
Grand Company manufactures and sells the following three products:

Economy Standard Deluxe
Unit sales 10,000 6,000 4,000
Unit sales price $50 $58 $70
Unit variable cost $30 $32 $36

Assume that total fixed cost is $344,400.

a. Compute the net income before income tax based on the sales volumes shown above.

Economy Standard Deluxe
Unit contribution margin Answer Answer Answer
Total contribution margin Answer Answer Answer

Net income before income tax: $Answer

b. Compute the break-even point in total dollars of revenue and in unit sales volume for each product.
Enter product mix answers in decimal form.

Product Product Mix Contribution Margin per unit Weighted average unit contribution margin
Economy Answer Answer Answer
Standard Answer Answer Answer
Deluxe Answer Answer Answer
Answer

Break-even Answer units   

Product Break -even Units Unit Sales Price Break-even Sales Revenue
Economy Answer Answer Answer
Standard Answer Answer Answer
Deluxe Answer Answer Answer
Answer

c. Prove your break-even calculations by computing the total contribution margin related to your answer in requirement (b).

Product Break -even Units Unit Contribution Margin Total Contribution Margin
Economy Answer Answer Answer
Standard Answer Answer Answer
Deluxe Answer Answer Answer
Answer

PreviousSave AnswersNext
Answer #1

A. Calculation of net income before income tax

Economy Standard Deluxe
Unit contribution margin 20 26 34
Total contribution margin $200,000 156,000 136,000

Unit contribution margin= selling price per unit - variable cost per unit

Economy= 50-30=20

Standard=58-32=26

Deluxe=70-36=34.

Total contribution margin= unit sales × contribution margin per unit

Economy=10,000×20= 200,000

Standard=6,000×26= 156,000

Deluxe= 4,000×34= 136,000.

Net income before tax= total contribution margin- total fixed cost

= 492,000- 344,400= $147,600.

B.

Product Product mix contribution margin per unit weighted average contribution margin
Economy 0.5 20 10
Standard 0.3 26 7.8
Deluxe 0.2 34 6.8
24.6

Break even answer units

Break even units unit sales price break even sales revenue
Economy 7,000 50 350,000
Standard 4,200 58 243,600
Deluxe 2,800 70 196,000
789,600

Break even units =( fixed cost ÷total weighted average contribution margin)×weighted average contribution margin

Economy=( 344,400÷24.6)×10= 14,000

Standard= (344,400÷24.6)×7.8= 4,200

Deluxe= (344,400÷24.6)×6.8= 2,800.

C. Computation of total contribution margin

Product Break even units unit contribution margin total contribution margin
Economy 7,000 20 140,000
Standard 4,200 26 109,200
Deluxe 2,800 34 95,200
344,400

____×____

All the best

Any help comment below